OpenStudy (anonymous):

How do you find cos(11pi/6)?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you need a good trig cheat sheet !

OpenStudy (michele_laino):

hint: we have: \[\Large \frac{{11\pi }}{6} = 2\pi - \frac{\pi }{6}\]

OpenStudy (michele_laino):

so we can write: \[\Large \cos \left( {\frac{{11\pi }}{6}} \right) = \cos \left( {2\pi - \frac{\pi }{6}} \right)\]

OpenStudy (misty1212):

find \(\frac{11\pi}{6}\) on the unit circle on the the last page of the cheat sheet, then look at the coordinates of the corresponding point the first coordinate is cosine

OpenStudy (anonymous):

Thanks guys

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

What if you need to find tangent?

OpenStudy (michele_laino):

you can apply this identity: \[\Large \begin{gathered} \tan \left( {\frac{{11\pi }}{6}} \right) = \tan \left( {2\pi - \frac{\pi }{6}} \right) = \hfill \\ \hfill \\ = \frac{{\tan \left( {2\pi } \right) - \tan \left( {\frac{\pi }{6}} \right)}}{{1 + \tan \left( {2\pi } \right)\tan \left( {\frac{\pi }{6}} \right)}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

How do you know tan(11pi/6) equals tan(2pi-pi/6)?